# A Deep Learning Framework for Disentangling Triangle Singularity and Pole-Based Enhancements

### v2

@@ IV.3 Training and testing
Figure 7: Training performance of DNN models using optimizers AdaGrad and AMSGrad with different hidden-layer architectures indicated above each plot. The blue lines are for the training dataset and the orange lines are for the testing set.
Figure 8: Training performance of DNN models using optimizer SMORMS3 with different hidden-layer architectures indicated above each plot. The blue lines are for the training dataset and the orange lines are for the testing set. We select the leftmost one as our DNN A and the rightmost as DNN B for validation and inference.
@@
During training, the prepared dataset is fed into the DNN using some mini-batch procedure to add stochasticity in estimating the cost-function. It must be noted that we independently produced a separate set of “testing” data that is hidden from the DNN during training. The testing dataset consist of 320 line shapes for each of the classifications, constructed using completely different sets of parameter values from the training dataset. This is done to ensure that the DNN is not simply “memorizing” the training data and has the ability to generalize beyond them. The testing data are only fed into the DNN after every training loop to measure the accuracy of the trained model. During every loop, the weights and biases of the nodes are being optimized following some framework for gradient descent. For this, we use different optimizers available in Chainer Cha and mix them with our chosen architectures. Table 4 shows the nine optimizers we tested in our study. This entire training-optimizing-testing procedure is repeated across 1000 epochs. We find that some models learn the classification problem but are eventually reduced to random models after a few more training epochs. There are also models that do not learn the problem at all. This is evidence that the ambiguity problem among the line shapes is indeed worth investigating. It is, therefore, important to find the most optimal model for line shape analysis purposes. Figure 7 shows some of the models considered in this study that can learn the classification problem. Figure 8 shows that the SMORMS3 (squared mean over root mean squared cubed) gives the best performance both in accuracy and stability among the optimizers considered. This particular optimizer works well when dealing with gradients of high variance, which is applicable to DNNs with multiple hidden layers SMO ; Funk (2015).
We choose two optimal models for later validation and for drawing inferences from the experimental data. The first one is the DNN model A with architecture 150-[250-100]-4 using SMORMS3 optimizer with final training accuracy of 98.0025% and testing of 96.0938%. The second one is the DNN model B with architecture 150-[250-250-250]-4 using SMORMS3 optimizer with a final training accuracy of 98.0175% and testing of 93.9062%. We have specifically chosen these two since one has the highest final training accuracy and the other has the highest final testing accuracy among the trained DNN models. Accuracy is a measurement of the proportion of correct predictions made by the DNN over a certain dataset. It is expected to achieve a high training accuracy for a well-trained DNN indicating that the model is capable of learning the classification problem using the training dataset. However, a high testing accuracy is arguably more crucial since this implies that the DNN model is able to generalize beyond the training dataset. Our ideal trained DNN should be able to generate good predictions for the separate testing dataset not included during the training without suffering under or overfitting. In our chosen models, both the final training and testing accuracies are greater than 90% in which they satisfy industry standards for well-trained models. It must be noted that the divergence between the training and testing accuracies is not a sign of overfitting, but rather due to saturation. It is expected that the predictions made for the testing dataset may not always be as accurate compared to that for the training. What matters is that the accuracies did not significantly drop and the final accuracies are stable and maintained across multiple epochs. With this, we may now proceed to the validation and inference stages for the two trained DNN models.
@@ V.1 Validation of trained DNN models
Figure 9: Confusion matrices for DNN A with architecture 150-[250-100]-4 (a) and DNN B with 150-[250-250-250]-4 (b). Both DNN models utilized the SMORMS3 optimizer in the training.
For a more thorough assessment of a trained DNN model, one may construct and analyze its confusion matrix. This allows us to observe how well the DNN makes predictions for each of the classifications and analyze where it may encounter difficulty. We independently generate a separate “validation” dataset using different sets of parameters, similar to how we prepared the testing dataset. The validation dataset also consists of 320 line shapes for each classification. We feed these into the trained DNN models and match the true (or expected) classification assigned to the line shapes with their predicted classification from the DNN. Figure 9 presents the two confusion matrices for DNN A and DNN B which tabulates the number of line shapes for each true label vs. predicted label. The vertical columns are for the true labels while the horizontal rows are for the predicted labels. Note that the matrices constructed are not $N\times N$-dimensional since we separated the T00A and T00B datasets to further investigate how the DNN classifies the two. We note that both T00A and T00B datasets produced from triangle mechanisms must share the same correct triangle singularity classification with label (0). The overall label assignment capability of our trained DNN is generally satisfactory as revealed in the large contribution in the diagonal of the matrices, where the correct predictions are located. We can compute the accuracy of the DNN models by taking the sum of all correctly predicted line shapes divided by the total number of data. We find that DNN A has an accuracy of 96.4063%, while DNN B has an accuracy of 93.9844% based on its performance on the validation dataset. These accuracies closely match the testing accuracies obtained during DNN training in section IV.
The DNN models generally perform well with classifying the pole-based enhancements, as observed in the yellow tiles in the confusion matrices. This is an ideal result for our objective since our goal is to discriminate kinematical effects from true unstable state enhancements. The confusion matrices show that our DNN models are capable of correctly recognizing pole-based enhancements and do not confuse them with kinematical effects. Therefore, it shall be able to give confident predictions of unstable state enhancements when it recognizes one from the experimental data. Furthermore, it is important to note the small mismatch in the case of the triangle singularities. The small mismatches found in the last row of the matrices suggest that the triangle mechanism may sometimes be confused as label (3), which is the case of having one pole each in the $[bt]$ and $[bb]$-sheets. Recall that this pole configuration suggests a compact state. This result strongly supports our original problem concerning ambiguity between the line shapes of a triangle singularity and a hadronic resonance. Furthermore, T00B has a larger number of mismatches compared to T00A in both confusion matrices. As previously discussed, the triangle mechanism in Figure 3(b) can sometimes produce similar line shapes as that of the direct coupled-channel reaction in Figure 3(d). In the latter mechanism, a Breit-Wigner set of poles in the second and third Riemann sheets are displaced due to strong coupling to the higher mass channel. This observation implies that a similar triangle-like shape may be produced by having one pole each in [$bt$] and [$bb$] with slightly different real and imaginary energy pole values.
@@
In summary, the confusion matrices reflected the ambiguity concerning the line shapes produced by the triangle mechanisms. Nonetheless, the analysis showed that the trained DNN models have no problem recognizing and classifying pole-based enhancements. Considering this knowledge and the fact that both DNN models obtained high accuracies during the training, testing, and validation stages, we are confident in treating them as well-trained models fit for inference. It, therefore, demands that we apply the trained DNNs to perform an interpretation of the $P_{\psi}^{N}(4312)^{+}$ experimental data.
@@ V.2 Interpreting $P_{\psi}^{N}(4312)^{+}$
Figure 10: Interpretation of $P_{\psi}^{N}(4312)^{+}$ state by trained DNN models. Experimental line shapes were generated using (a) uniform distribution and (b) normal distribution from given experimental uncertainties.
@@
Recall that the design of our DNN takes in random energy points and the corresponding intensity values, in order to take advantage of the limited energy resolution and statistical uncertainties from the experimental data. It is expected that, due to the presence of error bars, the interpretation of the experimental data can also be spread out. This possible variation in the interpretation should be reflected in the final inference stage. We generated 3,000 line shapes from the experimental data by randomly selecting a representative point from each energy bin, using a uniform distribution. For the intensity axis, we adhere to the error bars present in the experimental data and similarly, randomly select a representative point for each intensity corresponding to our selected energy points. For this task, we used a uniform distribution and a normal (or Gaussian) distribution. For the normal distribution, we treat one-half of the error bar to be equal to $1\sigma$, as opposed to the standard $3\sigma$. This is done since we observed that the latter setup will yield final results similar to using a uniform distribution. Thus, for variety, we use the $1\sigma$ distribution. We feed the experimental line shapes into our two trained DNN models and obtain their predictions. The results are shown in Figure 10.
@@
Notice that the triangle singularity interpretation, using label (0), for the $P_{\psi}^{N}(4312)^{+}$ is consistently ruled out by the two trained DNN models. Although in the case of the DNN A inference using a normal distribution, there is a small count of predictions for label (0), as seen in Figure 10(b), it is only around 3% of the total inputted experimental line shapes and may be negligible. Recall that in the initial analysis done by LHCb in Aaij et al. (2019), the triangle interpretation was ruled out due to the lack of hadron-rescattering threshold and the unrealistic parameters required to reproduce the line shape, as discussed in section II. Here, in our result, the triangle interpretation is ruled out by using only pure line-shape analysis with the help of trained DNN. This result is made more significant when we recall that we purposefully used the unrealistic triangle mechanism in order to achieve line shapes with prominent peaks that fit the experimental data, yet it is still not favored by the DNN and may consequently be ruled out as an interpretation for the origin of the $P_{\psi}^{N}(4312)^{+}$ state.
@@
Furthermore, it is interesting to point out that the trained DNN favored the pole-based interpretation, with a large percentage of predictions falling under label (1). That is, the peak found below the $\Sigma_{c}^{+}\bar{D}^{0}$ threshold is most likely produced by an isolated pole in the second Riemann sheet. This result conforms with the expected line shape where the peak appears below the threshold when the pole is in the $[bt]$-sheet. Moreover, recall our observation in section III that pole-based enhancements generally have a narrower width (FWHM) compared to the triangle and can fit the experimental uncertainties better. This may be one of the reasons why it is preferred by the DNN. Thus, naively speaking, a hadronic molecule interpretation corresponding to the label (1) classification may be inferred from our result as being the most favorable, which is indeed the case being argued by multiple studies Chen et al. (2019); Cheng and Liu (2019); Du et al. (2020). However, it has already been shown in Santos and Sombillo (2023) that such line shape is prone to ambiguous pole structure since it can also be produced by three poles distributed in each unphysical Riemann sheet. Addressing this ambiguity requires careful treatment where the contribution from the off-diagonal elements of the $S$-matrix must be included. Such a task is currently beyond the scope of the present study.
@@
Lastly, we note that even if two different probability distributions were used to produce the experimental line shapes used for inference, the results in both cases remain consistent. This demonstrates that the inference drawn from the trained DNN is robust and does not depend on how the inference dataset is generated from the experimental data. Therefore, to reiterate, using our deep learning framework, we have reached a similar conclusion with previous analysis that the triangle singularity may in fact be confidently ruled out for the $P_{\psi}^{N}(4312)^{+}$ state.
@@ VI Conclusion and Outlook
In this study, we have demonstrated for the first time how a triangle singularity can be discriminated from pole-based enhancements through a deep learning framework. Using the formalism of triangle loops and uniformized poles, we produced training datasets that teach the DNN model to recognize line shapes arising from both kinematic and dynamic processes of a reaction. The trained DNN models were shown to perform at high accuracies and provide robust and consistent inferences on experimental data. Following this, we have effectively shown that the single triangle interpretation for the $P_{\psi}^{N}(4312)^{+}$ hidden-charm pentaquark state may indeed be confidently ruled out solely based on a pure line-shape analysis. Our results properly reflect the argument of the LHCb that the parameters of the triangle mechanism are unrealistic and are therefore an unlikely explanation of the observed signal Aaij et al. (2019). We must note that our result only rules out the single triangle singularity picture, but do not rule out the possible mixing of TS contribution with other dynamic mechanisms. There is also the more rigorous treatment of double triangle singularity that can produce similar structures Nakamura (2021).
@@
Moving forward, we aim to perform similar studies that cover a wider energy range if not the whole invariant mass distribution of the $J/\psi p$ and apply our framework on the $P_{c}(4440)^{+}$ and $P_{c}(4457)^{+}$ states. This analysis would require working beyond two-channel treatments and investigate the possible coupling of other channels and their corresponding mechanisms Yan et al. (2022); Yang et al. (2022); Yan et al. (2023). This would also provide us the opportunity to use realistic triangle pictures in our DNN and obtain valuable inferences on the observed signals. Similar studies may also be done to evaluate the triangle singularity interpretation for various other potential candidates Guo et al. (2020). Moreover, we recommend that further studies be done to establish a more definitive interpretation for the $P_{\psi}^{N}(4312)^{+}$ state such as accommodating more sophisticated pole configurations, probing even distant Riemann sheets, and settling the ambiguity problem revealed in Santos and Sombillo (2023).
@@
In closing, we would like to remark that our study is the first-ever known case of the development of a model-selection framework in hadron physics. The concept of model selection is mostly used in the fields of statistics and machine learning, wherein the most appropriate model is often selected among a set of candidates following some measure or criteria. This is done to obtain the best model that accurately explains experimental data while providing reliable predictions Zucchini (2000). In our case, we have developed a framework that can select the most favorable mechanism to interpret a measured enhancement, while eliminating those misleading ones. For instance, the pole-based interpretation was selected as the most likely origin of the $P_{\psi}^{N}(4312)^{+}$ signal by the trained DNN. Consequently, the triangle singularity interpretation was then ruled out for this state. It must be noted that model selection is only suggestive in nature and never completely confirmatory Ding et al. (2018). Hence, it still demands supporting studies that can provide in-depth analyses and reasonable conclusions. Although our model-selection framework is currently limited to the chosen mechanisms used to train the DNN, its effective demonstration proves that it can serve as a valuable tool in hadron physics for classifying exotic hadron candidates. The novel framework can be extended to include more theoretical models and potentials used in hadron spectroscopy. Exotic hadrons may then be analyzed using this framework in order to gain intuition on the origin of their signals.

### v3

@@ IV.3 Training and testing
Figure 7: Training performance of DNN models in Table 4. The blue lines are for the training dataset and the orange lines are for the testing set.
@@
During training, the prepared dataset is fed into the DNN using some mini-batch procedure to add stochasticity in estimating the cost-function. It must be noted that we independently produced a separate set of “testing” data that is hidden from the DNN during training. The testing dataset consists of 320 line shapes for each classification, constructed using completely different sets of parameter values from the training dataset. This is done to ensure that the DNN is not simply “memorizing” and to check if the model can generalize beyond the training set. The testing data are only fed into the DNN after every training epoch to measure the accuracy of the trained model. During every loop, the weights and biases of the nodes are being optimized following some framework for gradient descent. For this, we use different optimizers available in Chainer Cha and mix them with our chosen architectures. Different optimizers are used such as Adam, AdaDelta, AdaGrad,AMSGrad, AdaBound, AMSBound, RMSprop, RMSpropGraves, and SMORMS3. We find that some models learn the classification problem but are eventually reduced to random models after a few more training epochs. There are also models that do not learn the problem at all. This is evidence that disentangling the nature of enhancements is a non-trivial task. For the different combinations of chosen architecture, only the AdaGrad, AMSGrad, and SMORMS3 show promising results. We label our chosen models according to Table 4 for reference.
Fig. 7 illustrates the various DNN models capable of learning the classification problem. The SMORMS3 (squared mean over root mean squared cubed) optimizer demonstrates rapid convergence, achieving higher training and testing accuracy compared to AdaGrad and AMSGrad. SMORMS3 is particularly effective in handling gradients with high variance, making it suitable for DNNs with multiple hidden layers SMO ; Funk (2015). While SMORMS3 shows a divergence between testing and training results, the primary focus remains on achieving high accuracy, which SMORMS3 consistently delivers. Additionally, other optimizers such as AdaGrad and AMSGrad also achieved high accuracy, further emphasizing the robustness of these methods.
Estimating the uncertainty of the reported accuracy, ideally without additional computational cost during training, is crucial. This can be achieved using the snapshot ensemble method, which builds an ensemble of models from different epoch states of the DNN Huang et al. (2017). Following this principle, we trained nine models for up to 1,050 epochs. We then used the accuracy values from epochs 950 to 1,050 to calculate the mean and standard deviation. The standard deviation of the 100 epoch states of a particular DNN model provides our estimation of the uncertainty in the accuracy. The results are presented in Table 5. In addition to the estimation of uncertainty in the accuracy, the snapshot ensemble method will be used in the final inference stage.
DNN Model Training Testing
Mean SD Mean SD
DNN 1 0.829 0.056 0.811 0.062
DNN 2 0.876 0.046 0.863 0.052
DNN 3 0.890 0.044 0.866 0.063
DNN 4 0.945 0.016 0.920 0.028
DNN 5 0.943 0.023 0.915 0.036
DNN 6 0.944 0.021 0.922 0.040
DNN 7 0.975 0.004 0.937 0.003
DNN 8 0.971 0.005 0.944 0.005
DNN 9 0.974 0.002 0.940 0.013
Table 5: Training and Testing Accuracy with estimated mean and uncertainty (SD) using the snapshot ensemble method for each DNN model listed in Table 4.
@@ V.1 Validation of trained DNN models
Figure 8: Confusion matrices for the nine DNN models. T00A, T00B, P01, P02, and P03 are the true labels while 0,1,2, and 3 are the predicted labels. The architectures and optimizers used are shown in Table 4.
During training, the testing accuracy plateaued below the training accuracy, indicating that some classifications in our training dataset are difficult to distinguish. To identify these confusing line shapes, we calculated the confusion matrix, focusing on the state of each DNN at epoch 1000. For validation, we generated an independent set of 320 datasets per classification. It is also useful to separate the two triangle enhancements considered in the study to determine which mechanism is most likely confused with the pole-based structure. Note that both T00A and T00B datasets produced from triangle mechanisms must fall under the label (0) to be considered correct. Fig. 8 shows the confusion matrices for the different DNN models, with rows corresponding to the true labels and columns to the predictions. Notice that the triangle mechanism in Fig. 3(a) (label T00A) is less likely to be confused with the pole-shadow pole configuration (label 3) compared to the other triangle. One possible reason for this distinction is the presence of cusp around the $\Lambda_{c}^{+}\bar{D}^{*0}$ threshold due to the final state interaction, despite being small, might be aiding the DNN in the classification. On the other hand, the second triangle diagram has no such cusp structure within the energy region considered in the analysis since $J/\psi p$ is already the lowest mass channel. Consequently, the triangle mechanism in Fig. 3(b) (label T00B) will most likely resemble a Breit-Wigner like line shape (label 3) as is evident in the confusion matrices. If the final inference using the experimental data favor either the label (0) or (3), then further analysis must be done.
@@
In general, The DNN models perform well with classifying the pole-based enhancements, as observed in the yellow tiles in the confusion matrices. This is an ideal result since our goal is to discriminate kinematical effects from dynamic pole structures. The confusion matrices show that our DNN models are capable of correctly recognizing the dynamic pole structures and will not mislabel them as triangle enhancements. To summarize, the confusion matrices reflected the ambiguity concerning the line shapes produced by the triangle mechanisms. Despite this, the analysis showed that the trained DNN models have no problem recognizing and classifying pole-based enhancements. Considering this and the fact that all trained DNN models obtained high accuracies during the training, testing, and validation stages, we move forward in applying them to perform an interpretation of the $P_{\psi}^{N}(4312)^{+}$ experimental data.
@@ V.2 Interpreting $P_{\psi}^{N}(4312)^{+}$
Table 6: The final inference result using the LHCb data. The mean count out of 3000 bootstrapped inference amplitudes and their corresponding standard deviations obtained by the snapshot ensemble approach.
Triangle 1 pole in $[bt]$ 1 pole in $[tb]$ 1 pole each in $[bt]$ and $[bb]$
DNN 1 $261\pm 39.4$ $1930\pm 166$ $0.248\pm 0.454$ $804\pm 169$
DNN 2 $18.4\pm 4.34$ $1998\pm 97.5$ $969\pm 93.6$ $14.5\pm 7.23$
DNN 3 $0.653\pm 1.29$ $2960\pm 22.2$ $0\pm 0$ $36.0\pm 21.7$
DNN 4 $0\pm 0$ $2990\pm 79.4$ $0\pm 0$ $14.2\pm 79.4$
DNN 5 $0\pm 0$ $2660\pm 274$ $0\pm 0$ $337\pm 274$
DNN 6 $0.436\pm 3.77$ $2999\pm 9.73$ $0\pm 0$ $0.891\pm 8.91$
DNN 7 $1\pm 0$ $2590\pm 16.3$ $0\pm 0$ $414\pm 16.3$
DNN 8 $0\pm 0$ $2999\pm 0.394$ $0\pm 0$ $1.06\pm 0.394$
DNN 9 $0\pm 0$ $2870\pm 14.3$ $0\pm 0$ $127\pm 14.3$
@@
Recall that the design of our DNN takes in random energy points and the corresponding intensity values, in order to take advantage of the energy resolution and statistical uncertainties from the experimental data. It is expected that, due to the presence of error bars, the interpretation of the experimental data can also be spread out. This possible variation in the interpretation should be reflected in the final inference stage. We performed bootstrapping technique and generated 3,000 line shapes from the experimental data by randomly selecting a representative point from each energy bin, using a uniform distribution. For the intensity axis, we exploit the error bars present in the experimental data and similarly, randomly select a representative point for each intensity corresponding to our selected energy points. We used snapshots of each DNN model to interpret the generated line shapes from the experimental data and calculated the mean and standard deviations of the predictions for each model. The standard deviation represents the confidence of the models in the interpretation as represented by the mean. The inference result is shown in Table 6.
@@
Notice that the triangle singularity interpretation for the $P_{\psi}^{N}(4312)^{+}$ is consistently ruled out by all of the trained DNN models. Although in the case of the DNN 1 inference, there is a small count of predictions for the triangle singularity, it is only around 8% of the total inputted experimental line shapes and has a significant deviation from pole-based interpretation. The remaining trained DNN models interprets the same but with negligible number of predictions corresponding to triangle singularity. Recall that in the initial analysis done by LHCb Aaij et al. (2019), the triangle interpretation was ruled out due to some physical constraints, as discussed in section II. Here, in our result, the triangle interpretation is ruled out by using only pure line-shape analysis with the help of trained DNN. This conclusion becomes even more significant when considering that we deliberately incorporated hypothetical triangle mechanisms in order to achieve line shapes with prominent peaks that fit the experimental data. Despite this, the DNN still does not favor the triangle interpretation, allowing it to be ruled out as a viable explanation for the origin of $P_{\psi}^{N}(4312)^{+}$.
@@
Furthermore, it is interesting to point out that the trained DNN models favor the pole-based interpretation, with a large percentage of predictions falling under the configuration of one pole in the [$bt$]-sheet. That is, the peak found below the $\Sigma_{c}^{+}\bar{D}^{0}$ threshold is most likely produced by an isolated pole in the second Riemann sheet. This result conforms with the expected line shape where the peak appears below the threshold when the pole is in the $[bt]$-sheet. Moreover, recall our observation in section III that pole-based enhancements generally have a narrower and asymmetric width compared to the triangle. This is one of the possible reasons why it is preferred by the DNN. Thus, naively speaking, a hadronic molecule interpretation corresponding to the label (1) classification may be inferred from our result as being the most favorable, which is indeed the case being argued by multiple studies Chen et al. (2019); Cheng and Liu (2019); Du et al. (2020). However, it has already been shown in Santos and Sombillo (2023) that such line shape is prone to ambiguous pole structure since it can also be produced by three poles distributed in each unphysical Riemann sheet. Addressing this ambiguity requires careful treatment where the contribution from the off-diagonal elements of the $S$-matrix must be included. A more rigorous probing of the pole structure of $P_{\psi}^{N}(4312)^{+}$ was recently addressed in Santos et al. (2024).
@@
Lastly, we also note that different probability distributions were employed in the bootstrapping process to generate the experimental line shapes used for inference. Remarkably, the results remained consistent across all cases. This demonstrates that the inference drawn from the trained DNN is robust and does not depend on how the inference dataset is generated from the experimental data. Therefore, to reiterate, our deep learning framework supports conclusions from previous analyses, indicating that the triangle singularity can indeed be ruled out as an explanation for the $P_{\psi}^{N}(4312)^{+}$ state.
@@ VI Conclusion and Outlook
In this study, we have demonstrated for the first time how a triangle singularity can be distinguished from pole-based enhancements through a deep learning framework. Using the formalism of triangle loops and uniformized poles, we produced training datasets that teach the DNN model to recognize line shapes arising from both kinematic and dynamic processes of a reaction. The trained DNN models were shown to provide reasonably high accuracy of discrimination between the two seemingly identical enhancements and also robust and consistent inference on the experimental data, despite the presence of experimental uncertainties. Using the trained DNN to analyze the line shape of the $J/\psi p$ invariant mass distribution near the $\Sigma_{c}^{+}\bar{D}^{0}$ threshold, the triangle singularity is consistently ruled out as the origin of the $P_{\psi}^{N}(4312)^{+}$ state.
The result of our analysis provides an independent validation to the analysis of LHCb that the triangle mechanism is an unlikely explanation of the enhancement just below the $\Sigma_{c}^{+}\bar{D}^{0}$ threshold Aaij et al. (2019). It must be pointed out that our present analysis does not claim that the pole-based enhancement for $P_{\psi}^{N}(4312)^{+}$ is the ultimate explanation for the observed signal. Other mechanisms should be included in the training dataset to make a more definitive description of the signal. Otherwise, with limited representation of possible mechanisms, the output of the DNN during the inference stage might be misleading. For example, our present analysis does not rule out the possibility that $P_{\psi}^{N}(4312)^{+}$ is due to a double triangle mechanism as discussed in Nakamura (2021). Inclusion of other mechanisms will be considered in our future investigations.
@@
Moving forward, we aim to perform similar studies that cover a wider energy range if not the whole invariant mass distribution of the $J/\psi p$ and apply our framework on the $P_{c}(4440)^{+}$ and $P_{c}(4457)^{+}$ states, which we have already partly accomplished in Co and Sombillo (2024). This analysis would require working beyond two-channel treatments and investigating the possible coupling of other channels and their corresponding mechanisms Yan et al. (2022); Yang et al. (2022); Yan et al. (2023). Mixing of different processes that can enhance possible suppressed triangle mechanisms should also be considered so that the parameters used in the training can still be constrained physically. Similar studies may also be done to evaluate the triangle singularity interpretation for various other potential candidates Guo et al. (2020). Moreover, we recommend that further studies be done to establish a more definitive interpretation for the $P_{\psi}^{N}(4312)^{+}$ state such as accommodating more sophisticated pole configurations, probing even distant Riemann sheets, and settling the ambiguity problem revealed in Santos and Sombillo (2023); Santos et al. (2024).
@@
In closing, we would like to remark that our study is the first-ever known case of the development of a model-selection framework in hadron physics. The concept of model selection is mostly used in the fields of statistics and machine learning, wherein the most appropriate model is often selected among a set of candidates following some measure or criteria. This is done to obtain the best model that accurately explains experimental data while providing reliable predictions Zucchini (2000). In our case, we have developed a framework that can select the most favorable mechanism to interpret a measured enhancement, while eliminating those misleading ones. For instance, the pole-based interpretation was selected as the most likely origin of the $P_{\psi}^{N}(4312)^{+}$ signal by the trained DNN. Consequently, the triangle singularity interpretation was then ruled out for this state. It must be noted that model selection is only suggestive in nature and never completely confirmatory Ding et al. (2018). Hence, it still demands supporting studies that can provide in-depth analyses and reasonable conclusions. Although our model-selection framework is currently limited to the chosen mechanisms used to train the DNN, its effective demonstration proves that it can serve as a valuable tool in hadron physics for classifying exotic hadron candidates. The novel framework offers an alternative approach to studying line shapes and supplements the standard fitting methods. It can be extended to include more theoretical models and potentials used in hadron spectroscopy. Exotic hadrons may then be analyzed using this framework in order to gain stronger intuition on the origin of their signals.